OpenStudy (anonymous):

20 Over 30 In simplest form? And 8 Over 40

OpenStudy (accessdenied):

\[\frac{20}{30} = \frac{2*10}{3*10} = \frac{2}{3} \] 2/3. Maybe that will help you figure out the second? (Find a common factor in the numerator and denominator and eliminate it)